Circular No. 7493

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2000dk IN NGC 382 S. Beckmann and W. D. Li, University of California at Berkeley, on

behalf of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), report the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of an apparent supernova on unfiltered images taken on Sept. 18.3 (mag ∼ 16.0) and 19.2 UT (mag ∼ 15.8). The new object is located at α = 1^{h}07^{m}23^{s}.52, δ = +32^{o}24′23′′.2 (equinox 2000.0), which is 5′′.3 west and 8′′.7 north of the nucleus of NGC 382. A KAIT image of the same field taken on Sept. 11.3 already showed a hint of SN 2000dk, while an image taken on Sept. 6.3 showed nothing at this position (limiting mag ∼ 19.0). NGC 382 is interacting with the galaxy NGC 383.

COMET C/1999 T1 (MCNAUGHT-HARTLEY) Ephemeris from orbital elements on MPC 41158:

2000TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Sept. 13 8^{h}50.^{m}39 −42^{o}49.^{′}9 2.132 1.816 58.^{o}2 28.^{o}1 10.2
23 9 30.45 −43 03.2 2.084 1.710 54.6 28.6 10.0
Oct. 3 10 10.00 −42 45.5 2.043 1.608 50.7 28.8  9.7
13 10 48.24 −41 55.5 2.007 1.511 46.8 28.8  9.4
23 11 24.69 −40 33.1 1.968 1.421 43.4 28.8  9.2
Nov. 2 11 59.13 −38 38.4 1.924 1.340 40.8 28.9  8.9
12 12 31.59 −36 10.3 1.871 1.273 39.4 29.6  8.7
22 13 02.36 −33 05.9 1.806 1.220 39.6 31.1  8.5
Dec. 2 13 31.82 −29 21.0 1.729 1.186 41.4 33.4  8.3
12 14 00.38 −24 48.5 1.642 1.172 44.9 36.3  8.1
22 14 28.48 −19 20.2 1.549 1.179 49.6 39.4  8.0

Visual m_{1} estimates: Aug. 10.81 UT, 11.0 (M. Mattiazzo, Wallaroo, S. Australia, 0.20-m reflector); Sept. 4.82, 10.8 (A. Pearce, Nedlands, W. Australia, 0.20-m reflector); 11.88, 10.2 (Pearce).

SUPERNOVA 2000cx IN NGC 524 Unfiltered CCD magnitude estimates: Aug. 2.09 UT, 13.5 (K. Sárnecz-

ky and L. Kiss, Szeged, Hungary); 26.960, 15.0 (E. Prosperi, Castelmartini, Italy); 29.951, 15.2 (Prosperi); Sept. 6.883, 15.6 (Prosperi); 9.997, 15.8 (Prosperi).

2000 September 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green